• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name brads

    Log entry time 13:45:06 on March 12, 2012

    Entry number 365983

    Followups:

    keyword=Initial plan for when beam returns -- Monday

    The TE is in progress and should complete around 2--2:30pm. BEFORE the Hall goes back to beam permit we need to install the sieves for both the LHRS and the RHRS.
    1. Sieve data runs with both arms at the production setting (2.228 GeV/c). Do NOT change any spectrometer or other magnet field settings. Take 30--45min of data for both arms, you will need to adjust the prescales to get the CODA rates back up. I presume the target nose will need to be drained of He for this too.
    2. Sieve data runs with both arms at 2.251 GeV/c). We already have this for the LHRS, so it is only the RHRS that is useful here. Doug does not need this for GEp -- we will only do this if it will be useful for g2p. Talk to the RC before
    3. Escorted Controlled Access to remove sieves from both arms
    4. Bleedthrough measurement using HRS singles as our metric. Doug will followup with a specific procedure.
    5. BCM calibration (Alex). 0 -- 200nA in 25nA steps -- let's get this done and analyzed ASAP, and update the coefficients everywhere so we can remove some of the beam current ambiguity that we've been dealing with.
    From this point there are two branches -- we will update as information comes in.
    1. Moller measurement. Priority for Swing if Sasha (Moller expert) and Yves (optics) OK it. (4 hours)
    2. Calorimeter BCM cross-calibration (Alex, 3 hours)
    3. Inelastic data when the above is done until Tuesday DAY. (More details on this to come. We will discuss it at the 4:15pm g2p meeting this afternoon.)


    A copy of this log entry has been emailed to: jpchen,slifer,camsonne,doug,rom,brads,jixie,rom,glamazdi